OpenStudy (anonymous):

Find the correct set up for finding the derivative of f(x)=√x using the definition of the derivative

myininaya (myininaya):

Definition of derivative of f is \[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

myininaya (myininaya):

\[f(x)=\sqrt{x} \text{ so } f(x+h)=?\]